uni much higher does it go? Use g-9.81 m/s2 Units [m/s
Answer #1

1) We will use the equation:

Delta x = ut - rac{1}{2}gt^2

Here 0.589772

and t = 0.183 s

Then 0.589 = u(0.183)-(9.81)(0.183)2

u-4.12m/s

2) Now we will use:

v = u - gt

v = 4.12 - (9.81)(0.183) = 2.325 m/s

3) At the maximum height v = 0 m/s

We will use:

v^2 = u^2 - 2gx, where x is the total height jumped.

иг 4.122 2(9.81)0.865m

Hence the armadillo goes = 0.865 - 0.589 = 0.276 higher than given.